OpenStudy (anonymous):

Given b = - 4 and h = -2, what is the equation of the graph if the parent function is y = square root of x?

OpenStudy (anonymous):

Also could you explain it i don't really understand what it is asking

OpenStudy (anonymous):

i believe, you are given 2 values, b and h it then asks you to use these values to create an equation it wants the equation to be in the form of y = sqrt(x)

OpenStudy (anonymous):

i believe the h corresponds to an x value, and b corresponds to a y value. it's basically asking you to do this; y = sqrt(x - h) - b where h and b are constants

OpenStudy (anonymous):

someone correct me if i'm wrong, cause i dont remember the general formula of a sqrt function

OpenStudy (anonymous):

it's kind've like y = mx + b if i told you i wanted an equation of a line, with b = -4 and m = 2, you could do that, right?

OpenStudy (anonymous):

same thing going on here, but instead of an equation for a line it's the equation for a square root

OpenStudy (anonymous):

and actually, the formula should be this: y = sqrt(x - h) + b not - b as i had written earlier if any of this has been confusing just ask

OpenStudy (anonymous):

Its a little confusing because it is giving me options to answer with such as y= sqrt4x+8 y= sqrt4x-8 y= sqrt-4x-8 y= sqrt-4x+8

OpenStudy (anonymous):

hmm... i'm not exactly sure what it is asking then are there any relevant equations listed somewhere in your book? such as a more general form for the sqrt function? i think the information that's given is useless, unless we know how b and h are used